OpenStudy (anonymous):

I will fan and medal!!! A rectangular prism is completely packed with 90 cubes of edge length fraction 1 over 2 inch, without any gap or overlap. Which of the following describes the volume of this rectangular prism? 11 unit cubes and 1 smaller cube of volume fraction 1 over 8 cubic inch 22 unit cubes and 4 smaller cubes of volume fraction 1 over 8 cubic inch each 11 unit cubes and 2 smaller cubes of volume fraction 1 over 8 cubic inch each 22 unit cubes and 2 smaller cubes of volume fraction 1 over 8 cubic inch each

OpenStudy (danjs):

The volume of each those small cubes is \[\large V = (\frac{ 1 }{ 2 })^3 = \frac{ 1 }{ 8 }\]

OpenStudy (danjs):

90 of those will have a total volume for the prism of; 90/8

OpenStudy (danjs):

A "unit cube" is just a volume of 1 each... The total Prism volume of 90/8 = 11 + 2/8

OpenStudy (danjs):

You would need 11 unit cubes and 2 of those small cubes , to total 90/8